# Second-order derivations of functions spaces — a characterization of second-order differential operators

### v2

@@ Abstract
for operators $D\colon\mathscr{C}^{k}(\Omega)\to\mathscr{C}(\Omega)$, where $k$ is a given nonnegative integer and the above identity is supposed to hold for all $f,g,h\in\mathscr{C}^{k}(\Omega)$. We show that besides the operators of first and second derivative, there are more solutions to this equation. Some special cases characterizing differential operators are also studied.
@@ 1 Introduction
for operators $T,A\colon\mathscr{C}^{k}(\Omega)\to\mathscr{C}(\Omega)$ In this case, the precise statement is more complicated. We will need two definitions before we can formulate their result.
@@
$\begin{array}[]{rcl}T(f)(x)&=&c(x)^{2}\cdot f^{\prime\prime}(x)+R(f)(x)\\[5.69054pt] A(f)(x)&=&c(x)\cdot f^{\prime}(x)\end{array}\qquad\left(f\in\mathscr{C}^{k}(\Omega),x\in\Omega\right),$
@@
$\begin{array}[]{rcl}T(f)(x)&=&c(x)^{2}\cdot f^{\prime\prime}(x)+b(x)\cdot f^{\prime}(x)\\[5.69054pt] A(f)(x)&=&c(x)\cdot f^{\prime}(x)\end{array}\qquad\left(f\in\mathscr{C}^{k}(\Omega),x\in\Omega\right).$
@@ 2 Operator relations and their connections
Thus the operator $A$ does not fulfill the requirements of Definition 2. So $A$ depends trivially on the derivative.
@@ 3 Results
Generalized monomials of degree 00 are constant functions, generalized monomials of degree $1$ are additive functions. Furthermore, generalized monomials of degree $2$ are called quadratic functions.
@@
Let $S$ be a commutative cancellative semigroup and $H$ be a commutative group. Let $G=S-S$ be a natural extension of $S$. Assume that multiplication by $m!$ is bijective on $H$. Let $f\colon S\to H$ be a solution of
@@
$D(f)(x)=c_{0}(x)f(x)\ln(f(x))+c_{1}(x)f^{\prime}(x)+c_{2}(x)f^{\prime\prime}(x)+d_{00}(x)f(x)\left(\ln(f(x))\right)^{2}$ (8)
holds for all $f\in\mathscr{C}^{k}(\Omega)$ and $x\in\Omega$. Further,
@@
If $D\colon\mathscr{C}^{k}(\Omega)\to\mathscr{C}(\Omega)$ satisfies equation (5), then due to Lemma 1 and Proposition 1, there exists a function $F\colon\Omega\times\mathbb{R}^{k+1}\to\mathbb{R}$ such that
@@
Since $D$ is pointwise localized, so does the operator $P$. Thus there exists a function $G\colon\mathbb{R}^{k+1}\times\Omega\to\mathscr{C}(\Omega)$ such that
@@
satisfies an Aichinger-type functional equation. At the same time, for every fixed $g\in\mathscr{C}^{k}(\Omega)$, the function $P(g)$ is continuous on $\Omega$. Thus for any $g\in\mathscr{C}^{k}(\Omega)$, the function $x\longmapsto G(x,g(x),\ldots,g^{(k)}(x))$ is continuous on $\Omega$, as well. Therefore, due to Corollary 2, there exist continuous functions $c\colon\Omega\to\mathbb{R}^{k+1}$, $d\colon\Omega\to\mathscr{M}_{(k+1)\times(k+1)}(\mathbb{R})$ such that
@@
If $f\in\mathscr{C}^{k}(\Omega)$ and $f(x)<0$ holds at a point $x\in\Omega$, then there exists an open interval $J\subset\Omega$ containing the point $x$ such that $f|_{J}<0$. Thus, we can find $g\in\mathscr{C}^{k}(\Omega)$ such that $g(x)<0$ for all $x\in\Omega$ and $f|_{J}=g|_{J}$. Thus Lemma 1 yields that $D(f)(x)=D(g)(x)$. Therefore, without loss of generality, we can assume that $f(x)<0$ for all $x\in\Omega$. Then $f=-|f|$ holds and if we substitute
@@
in (5), then due to Lemma 1, we obtain $D(-|f|)=-D(|f|)$. Thus we have
@@
At the same time, the domain of the operator $D$ is $\mathscr{C}^{k}(\Omega)$ yielding that the above formula should also hold for functions that have zeros in their range, too. Therefore, we will show that $c_{i}=0$ for $i\geq 3$ and $d_{ij}=0$ for $i+j>2$.
Due to Lemma 2, the order is singularity of the function $f\cdot(\ln\circ|f|)^{(i)}$ is $O\left(\frac{1}{f^{i-1}}\right)$.
@@
Since $\lim_{x\to 0}x\ln(|x|)=0$ and $\lim_{x\to 0}x\ln(|x|)^{2}=0$, in the above formula we adopt $00\ln(0)=$ 0 and $0\ln(0)^{2}=0$.
@@
Conversely, the operator $D$ given by the formula (10) annihilates all polynomials of degree at most $1$ and satisfies (5).
@@
A natural question is whether one can replace an operator relation containing more than one mapping by a relation with just one map. Similar question have been discussed recently in [4] and [5] in connection with other equations. It turns out that when we assume the linearity of the operator in question, then such weakening of the remaining assumption is possible. Before we proceed, we will need some auxiliary notions.
@@
and at the same time it is not a polynomial of order at most $2$. Such mappings do exist and can be chosen to be continuously differentiable, see e.g. the monograph of Kuczma [9, Lemma 12.1]. In fact, there exist a subinterval $I_{0}\subset]1,+\infty[$ such that every map $\varphi_{0}$ defined on $I_{0}$ can be extended to a solution of (12) defined on ]1,+[]1,$\infty$[] 1 , + ∞ [.
@@ Acknowledgment.
The research of E. Gselmann has been supported by project no. K134191 that has been implemented with the support provided by the National Research, Development and Innovation Fund of Hungary, financed under the K_20 funding scheme.
This paper was supported by the János Bolyai Research Scholarship of the Hungarian Academy of Sciences.

### v3

# Second-order derivations of function spaces — a characterization of second-order differential operators

@@ Abstract
for operators $D\colon\mathscr{C}^{k}(\Omega)\to\mathscr{C}(\Omega)$, where $k$ is a given nonnegative integer and the above identity is supposed to hold for all $f,g,h\in\mathscr{C}^{k}(\Omega)$. We show that besides the operators of first and second derivative, there are more solutions to this equation, and we characterize all solutions. Some special cases characterizing differential operators are also studied.
@@ 1 Introduction
for operators $T,A\colon\mathscr{C}^{k}(\Omega)\to\mathscr{C}(\Omega)$ . In this case, the precise statement is more complicated. We will need two definitions before we can formulate their result.
@@
$\begin{array}[]{rcl}T(f)(x)&=&\frac{1}{2}c(x)^{2}\cdot f^{\prime\prime}(x)+R(f)(x)\\[5.69054pt] A(f)(x)&=&c(x)\cdot f^{\prime}(x)\end{array}\qquad\left(f\in\mathscr{C}^{k}(\Omega),x\in\Omega\right),$
@@
$\begin{array}[]{rcl}T(f)(x)&=&\frac{1}{2}c(x)^{2}\cdot f^{\prime\prime}(x)+b(x)\cdot f^{\prime}(x)\\[5.69054pt] A(f)(x)&=&c(x)\cdot f^{\prime}(x)\end{array}\qquad\left(f\in\mathscr{C}^{k}(\Omega),x\in\Omega\right).$
@@ 2 Operator relations and their connections
Thus the operator $A$ does not fulfill the requirements of Definition 2, since $A$ depends trivially on the derivative.
@@ 3 Results
Generalized monomials of degree 0 are constant functions, generalized monomials of degree $1$ are additive functions. Furthermore, generalized monomials of degree $2$ are called quadratic functions.
In the following, we rely on several results from [2, 3], whose statements we reproduce here for completeness.
@@
Let $S$ be a commutative cancellative semigroup and $H$ be a commutative group. Let G=SSG=S-S be a natural extension of $S$. Assume that multiplication by $m!$ is bijective on $H$. Let $f\colon S\to H$ be a solution of
@@
$D(f)(x)=c_{0}(x)f(x)\ln(|f(x)|)+c_{1}(x)f^{\prime}(x)+c_{2}(x)f^{\prime\prime}(x)+d_{00}(x)f(x)\left(\ln(|f(x)|)\right)^{2}$ (8)
holds for all $x\in\Omega$ and $f\in\mathscr{C}^{k}(\Omega)$. Further,
@@
If $D\colon\mathscr{C}^{k}(\Omega)\to\mathscr{C}(\Omega)$ satisfies equation (5), then by Lemma 1 and Proposition 1, there exists a function $F\colon\Omega\times\mathbb{R}^{k+1}\to\mathbb{R}$ such that
@@
Since $D$ is pointwise localized, so is the operator $P$. Thus there exists a function $G\colon\mathbb{R}^{k+1}\times\Omega\to\mathscr{C}(\Omega)$ such that
@@
satisfies an Aichinger-type functional equation. At the same time, for every fixed $g\in\mathscr{C}^{k}(\Omega)$, the function $P(g)$ is continuous on $\Omega$. Thus for any $g\in\mathscr{C}^{k}(\Omega)$, the function $x\longmapsto G(x,g(x),\ldots,g^{(k)}(x))$ is continuous on $\Omega$, as well. Therefore, by Corollary 2, there exist continuous functions $c\colon\Omega\to\mathbb{R}^{k+1}$, $d\colon\Omega\to\mathscr{M}_{(k+1)\times(k+1)}(\mathbb{R})$ such that
@@
If $f\in\mathscr{C}^{k}(\Omega)$ and $f(x)<0$ holds at a point $x\in\Omega$, then there exists an open interval $J\subset\Omega$ containing the point $x$ such that $f|_{J}<0$. Thus, we can find $g\in\mathscr{C}^{k}(\Omega)$ such that $g(x)<0$ for all $x\in\Omega$ and $f|_{J}=g|_{J}$. Thus Lemma 1 yields that $D(f)(x)=D(g)(x)$. Therefore, without loss of generality, we can assume that $f(x)<0$ for all $x\in\Omega$. Then f=|f|f=-|f| holds and if we substitute
@@
in (5), then due to Lemma 1, we obtain D(|f|)=D(|f|)D(-|f|)=-D(|f|). Thus we have
@@
At the same time, the domain of the operator $D$ is $\mathscr{C}^{k}(\Omega)$ yielding that the above formula should also hold for functions that have zeros in their range, too. Using this, we will show that $c_{i}=0$ for $i\geq 3$ and $d_{ij}=0$ for $i+j>2$.
In view of Lemma 2, the order is singularity of the function $f\cdot(\ln\circ|f|)^{(i)}$ is $O\left(\frac{1}{f^{i-1}}\right)$.
@@
Since $\lim_{x\to 0}x\ln(|x|)=0$ and $\lim_{x\to 0}x\ln(|x|)^{2}=0$, in the above formula we adopt $0\ln(0)=0$ and $0\ln(0)^{2}=0$.
@@
Conversely, the operator $D$ given by the formula (10) annihilates all polynomials of degree at most 11 and satisfies (5).
@@
A natural question is whether one can replace an operator relation containing more than one mapping by a relation with just one map. Similar questions have been discussed recently in [4] and [5] in connection with other equations. It turns out that when we assume the linearity of the operator in question, then such weakening of the remaining assumption is possible. Before we proceed, we will need some auxiliary notions.
@@
and at the same time it is not a polynomial of order at most $2$. Such mappings do exist and can be chosen to be continuously differentiable, see e.g. the monograph of Kuczma [9, Lemma 12.1]. In fact, there exist a subinterval $I_{0}\subset]1,+\infty[$ such that every map $\varphi_{0}$ defined on $I_{0}$ can be extended to a solution of (12) defined on $1,+\infty[$.
@@ Acknowledgment.
The research of E. Gselmann was supported by the János Bolyai Research Scholarship of the Hungarian Academy of Sciences.